# Realistic Cost to Execute Practical Quantum Circuits using Direct Clifford+T Lattice Surgery Compilation

### v2

@@ Acknowledgements.
This work was completed as part of the Defense Advanced Research Projects Agency Quantum Benchmarking program. This manuscript has been authored by UT-Battelle, LLC, under contract DE-AC05-00OR22725 with the US Department of Energy (DOE). The US government retains and the publisher, by accepting the article for publication, acknowledges that the US government retains a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this manuscript, or allow others to do so, for US government purposes. DOE will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan (https://www.energy.gov/downloads/doe-public-access-plan).

### v4

@@ Acknowledgements.
We would like to thank Peter Selinger for many great conversations, and for his support with the Quipper implementation of the GSE algorithms considered here. We would also like to thank Scott Wesley for his help with LinguaQuanta. The work of T. LeBlond and R. Bennink was supported by the Defense Advanced Research Projects Agency (DARPA) Quantum Benchmarking (QB) and Underexplored Systems for Utility-Scale Quantum Computing (US2QC) programs under award numbers HR00112580540001 and HR0011261528. The work of C.  Dean was supported by the DARPA QB program under award number HR001122C0066 and was supported by the Air Force Office of Scientific Research, Air Force Material Command, USAF under award number FA9550-21-1-0041. The work of G. Watkins was supported by the DARPA QB program under award numbers HR00112230006 and HR001121S0026 and was supported by the QuantERA grant EQUIP through the Academy of Finland, decision number 352188. The views, opinions and/or findings expressed are those of the author(s) and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government. This manuscript has been authored by UT-Battelle, LLC, under contract DE-AC05-00OR22725 with the US Department of Energy (DOE). The US government retains and the publisher, by accepting the article for publication, acknowledges that the US government retains a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this manuscript, or allow others to do so, for US government purposes. DOE will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan (https://www.energy.gov/downloads/doe-public-access-plan). This research used resources of the Compute and Data Environment for Science (CADES) at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725.